a deposit of $420 earns interest rates of 8.2 percent in the first year and 10.2 percent in the second year. What would be second year future value?
Answer #1

Future value formula

PV= present value=420

I= interest rate = 8.2 for first year

N= number of years

FV= PV (1+I)^n

Fv = 420 (1+0.082)=454.44

Next interest rate=10.2

PV= 454.44

FV=454.44(1+.102)

FV= 500.79

Or

FV= PV(1+i) ×(1+j)

i= first year interest rate

j= second year interest rate

FV= 420(1+.08)×(1+0.102)=500.79

Both method is same but explained in details in first method . Future value calculated using compound interest formula so answer is fv=500.79